# Depth-Breadth Synergy in RLVR: Unlocking LLM Reasoning Gains with Adaptive Exploration

### v4

@@ 4.2 Main Results
4.3 Training Dynamics and Ablation Study
Pass@128 performance surpasses the base model, peaks quickly, and then declines.
We conduct RLVR experiments with rollout size 8/32 to compare our DARS (with $N^{\mathrm{max}}=32$), the training dynamics of Pass@128 performance during training is shown in Figure 7. Across all settings, Pass@128 surpasses the base model during training, but declines after peaking, indicating that over-training with RLVR harms Pass@128 performance. Notably, DARS (with $N^{\mathrm{max}}=32$) incurs substantially less inference cost than naively scaling the rollout size to n = 32. Despite this being an unfair comparison in terms of computational expenditure, our DARS not only attains the highest peak Pass@128 performance but also outperforms all other settings.
Figure 8: Training dynamics of Pass@32/Pass@128 and Pass@1 performance with different training steps of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
Depth Training with DARS Improve Pass@K Performance and Training Efficiency.
Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1. Figure 8 plots Pass@128 against Pass@1 under a variety of experimental settings. It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@128 than the other settings.
@@ 4.3 Training Dynamics and Ablation Study
Depth-Breadth are Complementary in RLVR.
We further illustrate that Depth and Breadth are two complementary dimensions in RLVR. As shown in Figure 9, we present the Pass@1Pass@K training-dynamics curves for the Breadth, Depth, and the two-dimensional synergy approach DARS-Breadth. The farther the Pass@1Pass@K curve deviates outward, the more powerful the method. Our DARS-Breadth curves lie on the outermost envelope: it not only achieves the best Pass@1, but also simultaneously lifts Pass@K. This demonstrates the complementary roles of Depth and Breadth in RLVR training.
Figure 9: Complementary improve of Depth and Breadth Synergy for Pass@1 and Pass@K (K=128) performance.
Ablation Study on Base Model.
We further illustrate the effectiveness of our DARS method on Llama-3.1-8B. The results are shown in Table 3. Our DARS-ET-Breadth achieves both Pass@1 and Pass@128 performance compared to other baselines, which further illustrates the effectiveness of our method.
@@
Complete Pass@K Accuracy Curve.
We show the complete Pass@K curve for Llama-3.1-8B, Qwen2.5-Math-1.5B, and Qwen2.5-Math-7B in Figure 10. The 3 chosen models of DARS are: DARS-Llama-ET-Breadth, DARS-1.5B-HW-Breadth, and DARS-7B-HW-Breadth. DARS models demonstrate a breakthrough in the reasoning boundaries of the base model, especially on the LLama-3.1-8B model, where the improvement in Pass@k is particularly significant.
@@ 5 Related Works
5.1 RL for LLM Reasoning
Reinforcement Learning (RL) is widely used in the post-training phase of LLMs. Previously, a reward model was employed to conduct preference comparisons among responses (Ouyang et al., 2022b). Direct Preference Optimization (Rafailov et al., 2023) is further proposed to simplify the training pipeline of RL for LLMs, which directly uses pair-wise preference data for model optimization. Recent studies have shown that RLVR can significantly improve the reasoning capabilities of LLMs in domains such as mathematics and programming. OpenAI’s o1 model (Jaech et al., 2024) represents a major milestone in LLM reasoning. Shortly thereafter, DeepSeek-R1 (Guo et al., 2025) demonstrated that RLVR can directly elicit a base model’s slow-thinking capability—a setting now termed zero-RL (Li et al., 2025). Subsequently, a series of Large Reasoning Models (LRM), including Kimi 1.5 (Team et al., 2025), Gemini-Think (DeepMind, 2024), and QwQ (Qwen, 2024) were successively introduced. These works have inspired a wave of follow-up research: Zeng et al. (2025) conducted extensive experimental exploration of zero-RL, while Luo et al. (2025) leveraged RLVR to further boost the performance of open-source models developed from DeepSeek-R1. Among these works, one widely adopted RLVR algorithm is GRPO (Shao et al., 2024), which builds upon PPO (Schulman et al., 2017) and attains substantial performance gains by sampling multiple responses to compute the group-relative advantage. GRPO has inspired a series of subsequent algorithms, including DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), and Dr. GRPO (Liu et al., 2025b). However, both GRPO and its variants tend to undervalue high-difficulty problems under cumulative advantage, which compromises the Pass@K performance of RLVR.
5.2 Analysis of RLVR
With the rapid advancement of RLVR and the proliferation of open-source LRMs, many studies have begun to analyze the RLVR pipeline and these open LRMs. Several studies (Liu et al., 2025a; Zhao et al., 2025; Shah et al., 2025) indicates that the self-reflect and self-critique behaviors observed after RLVR originates from the base model rather than the RL process. Dang et al. (2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses. Subsequently, Yue et al. (2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model. These studies have sparked widespread concern about the capability ceiling of RLVR, and consequently, the Pass@K metric has become a focal point for diagnosing and potentially transcending the intrinsic limits imposed by the base model (Liang et al., 2025). This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
@@ 6 Conclusion
In this work, we identify and systematically analyze two critical yet under-explored dimensions in Reinforcement Learning with Verifiable Reward (RLVR): Depth (the hardest problems a model can learn) and Breadth (the number of instances processed per iteration). We reveal that existing GRPO-based methods suffer from a fundamental limitation—their cumulative advantage calculation disproportionately weights medium-difficulty problems while neglecting hard ones, creating a bottleneck for Pass@K performance. To address this issue, we propose Difficulty-Adaptive Rollout Sampling (DARS) to reallocate compute to the hard problems, boosting Pass@K with low extra inference cost. Furthermore, we demonstrate that large-breadth training acts as an orthogonal dimension for RLVR optimization. Scaling batch size and employing full-batch updates significantly boost Pass@1 performance while sustaining high token-level entropy. By unifying these dimensions in our DARS-Breadth framework, we achieve simultaneous gains in both Pass@1 and Pass@K, confirming the synergistic relationship between depth and breadth in RLVR.
@@
Appendix A Derivation of Additional Rollouts $\Delta n_{j}$
@@
Appendix B Training and Evaluation Details
B.1 Setting Details
@@
Following LUFFY (Yan et al., 2025), we use the default subset and filter out generations that are longer than 8192 tokens and those that are verified wrong by Math-Verify  [^1] https://github.com/huggingface/Math-Verify, resulting in 45k question-solution pairs. For training Llama-3.1-8B, we use the train split of MATH dataset. Our training framework is derived from Verl (Sheng et al., 2024) pipeline, which is a flexible, high-performance reinforcement-learning framework built for training large language-model agents. With native PyTorch support and efficient distributed training, Verl lets researchers quickly prototype and scale RL algorithms like PPO on GPUs. All of our experiments are conducted on H200 GPUs. At present, the LLM of our experiment is the Qwen2.5-Math series.
@@
Appendix C More Experimental Results
C.1 Ablation Study on std-based Advantage Computation
@@
Figure 11: Comparison of our DARS on std-based advantage computation.
C.2 Depth and Breadth Synergy for Pass@1 and Pass@32
In Section 4.3, we show the training dynamics of Pass@128-Pass@1 for DARS and baseline methods. To further illustrate the effectiveness of DARS, we show the training dynamics of Pass@32-Pass@1 in Figure 12. Our DARS significantly improves the Pass@32 performance compared to other methods.
@@ C.2 Depth and Breadth Synergy for Pass@1 and Pass@32
C.3 Comparison of ET/HW Schedule in Breadth Scaling
@@
Appendix D Discussion and Future Work

### v5

@@ 4.2 Main Results
Figure 8: Depth and Breadth Synergy for Pass@1 and Pass@K (K=128) performance.
@@
4.3 Training Dynamics and Ablation Study
Pass@128 performance surpasses the base model, peaks quickly, and then declines. We conduct RLVR experiments with rollout size 8/32 to compare our DARS (with $N^{\mathrm{max}}=32$), the training dynamics of Pass@128 performance during training is shown in Figure 6. Across all settings, Pass@128 surpasses the base model during training, but declines after peaking, indicating that over-training with RLVR harms Pass@128 performance. Notably, DARS (with $N^{\mathrm{max}}=32$) incurs substantially less inference cost than naively scaling the rollout size to n = 32. Despite this being an unfair comparison in terms of computational expenditure, our DARS not only attains the highest peak Pass@128 performance but also outperforms all other settings.
@@ 4.3 Training Dynamics and Ablation Study
Depth Training with DARS Improve Pass@K Performance and Training Efficiency. Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1. Figure 7 plots Pass@128 against Pass@1 under a variety of experimental settings. It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@128 than the other settings.
Figure 10: Comparison of our DARS-HW and GRPO-Baseline on AIME 2024. We show the average response length and accuracy for Qwen2.5-Math-1.5B and Qwen2.5-Math-7B. After training with DARS, the models can achieve higher accuracy by utilizing longer thinking process.
Depth-Breadth are Complementary in RLVR. We show that Depth and Breadth are two complementary dimensions in RLVR. As shown in Figure 8, we present the Pass@1Pass@K training-dynamics curves for the Breadth, Depth, and the two-dimensional synergy approach DARS-Breadth. The farther the Pass@1Pass@K curve deviates outward, the more powerful the method. Our DARS-Breadth curves lie on the outermost envelope: it not only achieves the best Pass@1, but also simultaneously lifts Pass@K. This demonstrates the complementary roles of Depth and Breadth.
Ablation Study on Base Model. We further illustrate the effectiveness of DARS on Llama-3.1-8B. The results are shown in Table 3. Our DARS-ET-Breadth achieves both the highest Pass@1 and Pass@128 performance, which further illustrates the effectiveness of our method. DARS also consistently improves the performance of openPangu (Chen et al., 2025), as shown in Appendix E.
DARS Elicits Longer Thought. We tracked the response length dynamics during the training of DARS and GRPO. As shown in Figure 10, on the challenging AIME 2024 benchmark, DARS consistently produces longer reasoning traces than the baseline and has higher accuracy. These results reveal that our DARS stimulates the model to perform deeper thinking to solve hard problems.
Complete Pass@K Accuracy Curve. We show the complete Pass@K curve for Llama-3.1-8B, Qwen2.5-Math-1.5B, and Qwen2.5-Math-7B in Figure 9. The 3 chosen models of DARS are: DARS-Llama-ET-Breadth, DARS-1.5B-HW-Breadth, and DARS-7B-HW-Breadth. DARS models demonstrate a breakthrough in the reasoning boundaries of the base model, especially on the LLama-3.1-8B model, where the improvement in Pass@k is particularly significant.
@@ 5 Related Works
Reinforcement Learning (RL) is now standard in post-training LLMs. After early reward-model pipelines (Ouyang et al., 2022a), Direct Preference Optimization (Rafailov et al., 2023) streamlined training by exploiting pairwise preferences. RL with verifiable rewards (RLVR) has since pushed reasoning benchmarks in math and code, culminating in OpenAI’s o1 (Jaech et al., 2024) and the zero-RL breakthrough of DeepSeek-R1 (Guo et al., 2025). Follow-up Large Reasoning Models—Kimi 1.5 (Team et al., 2025), Gemini-Think (DeepMind, 2024), QwQ (Qwen, 2024), and studies like (Zeng et al., 2025; Luo et al., 2025) further validate RLVR. The leading algorithm, GRPO (Shao et al., 2024), extends PPO (Schulman et al., 2017) with group-relative advantages, inspiring DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), and Dr. GRPO (Liu et al., 2025b). Yet GRPO and its variants systematically undervalue hard problems, hurting Pass@K. More related works are shown in Appendix A.
@@ 6 Conclusion
In this work, we reveal that GRPO-based RLVR methods under-weight hard problems due to cumulative-advantage bias, capping Pass@K. Our DARS sampler cheaply re-allocates rollouts to these hard instances, while large-breadth training with full-batch updates raises Pass@1. The unified DARS-Breadth framework jointly lifts Pass@1 and Pass@K, proving depth and breadth are synergistic levers in RLVR.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here
@@
Appendix A More Related Works
With the rapid advancement of RLVR and the proliferation of open-source LRMs, many studies have begun to analyze the RLVR pipeline and these open LRMs. Several studies (Liu et al., 2025a; Zhao et al., 2025; Shah et al., 2025) indicates that the self-reflect and self-critique behaviors observed after RLVR originates from the base model rather than the RL process. (Dang et al., 2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses. Subsequently, (Yue et al., 2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model. These studies have sparked widespread concern about the capability ceiling of RLVR, and consequently, the Pass@K metric has become a focal point for diagnosing and potentially transcending the intrinsic limits imposed by the base model (Liang et al., 2025). This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
Appendix B Derivation of Additional Rollouts $\Delta n_{j}$
@@
Appendix C Mathematical Derivation: Group Cumulative Advantage and GRPO Gradient Norm
In this appendix, we provide a detailed mathematical derivation demonstrating the relationship between Group Cumulative Advantage and the gradient norm in GRPO. This derivation substantiates the claim that $\mathcal{A}_{\text{group}}$ serves as an effective indicator of the model’s attention to specific problems during training.
C.1 GRPO Gradient Formulation
The GRPO objective function and its gradient are given by (we can remove the clip operation for simplified analysis):
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\bigg],$ (11)
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\nabla_{\theta}\rho_{i,t}(\theta)\bigg]$
$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\frac{\nabla_{\theta}\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}\bigg]$
$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\bigg].$ (12)
C.2 Derivation of Gradient Norm Upper Bound
We now derive the upper bound relationship between the gradient norm and Group Cumulative Advantage.
The gradient norm of our adopted GRPO algorithm is shown as the following:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|=\left\|\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right]\right\|.$ (13)
Since the norm is a convex function and expectation is linear, by Jensen’s inequality:
$\left\|\mathbb{E}[X]\right\|\leq\mathbb{E}[\|X\|]$ (14)
Thus:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\left\|\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (15)
Applying the triangle inequality to the inner summation:
$\left\|\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}a_{i,t}\right\|\leq\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|a_{i,t}\right\|,$ (16)
where:
$a_{i,t}=\frac{1}{\sum_{i=1}^{G}|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t}).$ (17)
Therefore:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (18)
Then we have,
$\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|=|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|.$ (19)
Thus:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (20)
We further take a boundedness assumption. in policy optimization, we assume the gradient log-probabilities $\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|$ are bounded. Furthermore, it’s vital to notice that the importance ratio $\rho_{i,t}(\theta)$ are also bounded through the clip operation in GRPO algorithm.
Thus, there exists a constant $C>0$ such that:
$|\rho_{i,t}(\theta)|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\leq C.$ (21)
Therefore:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|\right].$ (22)
Noting that $|\hat{A}_{i}|$ is independent of $t$ for fixed $i$:
$\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|=\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|.$ (23)
Thus:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\right].$ (24)
In this paper, we define the Group Cumulative Advantage as:
$\mathcal{A}_{\text{group}}=\sum_{i=1}^{G}|\hat{A}_{i}|.$ (25)
By the weighted arithmetic mean inequality:
$\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\leq\frac{\sum_{i=1}^{G}|o_{i}|\cdot\sum_{i=1}^{G}|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}=\sum_{i=1}^{G}|\hat{A}_{i}|=\mathcal{A}_{\text{group}}.$ (26)
Therefore:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\mathcal{A}_{\text{group}}\right].$ (27)
C.3 Theoretical Implications
This derivation establishes that Group Cumulative Advantage $\mathcal{A}_{\text{group}}$ provides an upper bound for the expected gradient norm in GRPO. Consequently: Larger $\mathcal{A}_{\text{group}}$ values indicate stronger gradient signals for the corresponding problem. The training process allocates more ”attention” to problems with higher $\mathcal{A}_{\text{group}}$ values during parameter updates. Therefore, we consider $\mathcal{A}_{\text{group}}$ serves as a mathematically grounded indicator of problem importance in GRPO training dynamics.
This theoretical foundation validates the use of Group Cumulative Advantage as a meaningful metric for analyzing training behavior and problem prioritization in GRPO.
Appendix D Training and Evaluation Details
@@
Following LUFFY (Yan et al., 2025), we use the default subset and filter out generations that are longer than 8192 tokens and those that are verified wrong by Math-Verify  [^1] https://github.com/huggingface/Math-Verify, resulting in 45k question-solution pairs. For training Llama-3.1-8B, we use the train split of MATH dataset. Our training framework is derived from Verl (Sheng et al., 2024) pipeline, which is a flexible, high-performance reinforcement-learning framework built for training large language-model agents. With native PyTorch support and efficient distributed training, Verl lets researchers quickly prototype and scale RL algorithms like PPO on GPUs. Following Dr. GRPO (Liu et al., 2025b), we remove the KL loss and the length normalization in GRPO. All of our experiments are conducted on H200 GPUs. At present, the LLM of our experiment is the Qwen2.5-Math series.
Figure 11: Comparison of our DARS on std-based advantage computation.
@@
Appendix E More Experimental Results
E.1 Ablation Study on std-based Advantage Computation
@@
E.2 Depth and Breadth Synergy for Pass@1 and Pass@32
In Section 10, we show the training dynamics of Pass@128-Pass@1 for DARS and baseline methods. To further illustrate the effectiveness of DARS, we show the training dynamics of Pass@32-Pass@1 in Figure 13. Our DARS significantly improves the Pass@32 performance compared to other methods.
@@ E.2 Depth and Breadth Synergy for Pass@1 and Pass@32
E.3 Comparison of ET/HW Schedule in Breadth Scaling
@@
Figure 14: Heat map of Pass@128 of Qwen2.5-Math series in different temperatures.
E.4 Impact of Temperature
Some researches (Karan & Du, 2025; Qin et al., 2025; Ni et al., 2025) indicates that temperature matters in LLM reasoning. To further illustrate the performance improvement under different temperatures, we additionally add the above experiments. The results show that the improvement of our method is consistent over different temperatures. The results are shown in Figure 14.
E.5 Consistent Improvement During RL Process
To further show that our method consistently improve model performance, we calculated the mean of Pass@128 and Pass@32 for the last 3 checkpoints of each method, as shown in Table 4.
Table 4: Average performance of Pass@1/32/128 for the last 3 checkpoints during training.
Model Pass@1 Pass@32 Pass@128
Qwen2.5-Math-1.5B as the Base Model
GRPO-Baseline 48.8 74.7 80.8
Depth-Naive 49.5 74.4 80.6
Breadth-Naive 51.4 74.4 79.8
DARS-HW 49.5 75.7 81.9
DARS-HW-Breadth 52.4 76.4 82.1
Qwen2.5-Math-7B as the Base Model
GRPO-Baseline 55.1 76.9 81.8
Depth-Naive 56.4 76.7 80.9
Breadth-Naive 57.2 76.7 81.3
DARS-HW 56.8 78.8 83.4
DARS-HW-Breadth 58.3 79.1 83.7
E.6 Performance of Other Model
We further evaluate our method on Qwen2.5-7B-Instruct. Following (Liang et al., 2025), we change the training data to Math12k. The results are shown in Table 5.
We further evaluate DARS with the openPangu-7B architecture on Huawei Ascend NPU, to investigate the generalization ability of DARS on different models and hardware architectures. The experimental results are reported in Table 6, and DARS improves the performance of the openPangu model consistently on all datasets.
As the results show, our method still outperforms the baseline in both the Pass@1 and Pass@K metrics.
Table 5: Overall performance of Pass@1 (Avg@128) and Pass@128 of Qwen2.5-7B-Instruct.
Model AIME24 Math500 Olympiad AMC Minerva Avg@128 Pass@128
Qwen2.5-7B-Instruct 11.9 72.3 37.1 42.2 31.9 47.2 80.3
GRPO-baseline 14.2 74.8 37.6 43.4 33.4 48.6 78.8
DARS-HW-Breadth 15.6 76.5 38.4 44.7 34.6 49.6 82.3
Table 6: Overall performance of DARS with openPangu on Ascend NPU.
Model AIME24 Math500 Olympiad AMC Minerva Avg@32 Pass@32
openPangu-7B 36.99 82.71 48.82 51.61 37.0 57.54 79.92
openPangu-7B + DARS 60.58 93.95 67.96 81.06 43.4 72.56 86.59
E.7 Explanation of Significant Pass@k Gain for Llama-3
The stronger gains for Llama-3 likely stem from its base model characteristics. Recent studies (Gandhi et al., 2025; Wang et al., 2025) suggest that base models like Llama-3 lack mid-training, causing many valid responses to appear only at higher sampling budgets. As shown in the table below, Llama-3-8B-Base exhibits a much larger performance gap when increasing (K) from 8 to 32/64 compared to Qwen models as shown in Table 7. This indicates that Llama-3 has more correct solutions “hidden” deeper in the sampling space. Our DARS method, by focusing on hard examples, effectively uncovers these “deep” correct samples, leading to more substantial Pass@K improvements for Llama-3.
Table 7: Pass@K gap comparison of Qwen2.5-Math-1.5/7B and LLama3-8B-Base
Model Pass@8 Pass@32 - Pass@8 Pass@64 - Pass@8
Qwen2.5-Math-1.5B 56.5 12.4 17.1
Qwen2.5-Math-7B 63.5 10.5 14.4
Llama3-8B-Base 17.2 19.1 27.3
E.8 Thinking Length Dynamics
This section investigates how DARS influences the reasoning length of LLMs. We tracked the response length dynamics during the training of Qwen2.5-Math-1.5B and 7B models. Our experiments reveal two key observations: (1) The training process shows a clear trend of increasing generation length, as shown in Figure 15. (2) When evaluated on AIME 2024, models trained with DARS consistently produce longer reasoning traces than the baseline, as shown in Figure 16 (For steps in multiples of 100, we take the average of 128 sampling results. To increase the density of data points, we also sample 16 times for every intermediate 50-step interval and take the average). These results provide concrete evidence that our DARS method successfully stimulates the model to perform deeper and more thorough thinking.
Figure 15: Training dynamics of response length for GRPO and DARS.
Figure 16: Statistical results of response length on AIME 2024 for GRPO and DARS.
Appendix F Discussion and Future Work